An object is thrown upward from point A with a velocity of 40ms-1 calculate the time it takes to return to point A
1
Expert's answer
2019-03-25T09:07:59-0400

Let's first find the time an object takes to reach the maximum height from the kinematic equation:


"v=v_0 - gt,"

here, "v_0" is initial velocity of the object, "v = 0" is the velocity of the object at the maximum height, "g" is the acceleration due to gravity and "t" is the time the object takes to reach the maximum height.

Then, we get:


"t = \dfrac{v_0}{g} = \dfrac{40 \dfrac{m}{s}}{9.8 \dfrac{m}{s^2}} = 4.1 s."

Finally, we can find the total time the object takes to return to point A:


"t_{tot} = 2t =2 \cdot 4.1 s = 8.2 s."

Answer:

"t_{tot} = 8.2 s."